# New theoretical Fe II templates for bright quasars

### v1

@@ 3.2 Red-to-blue UV Fe II wings ratio
In our next plot, we concentrate on the global properties of the UV part of Fe II emission which is important for the proper modelling of the Mg II line. The Fe II emission under the Mg II line is never close to zero, even right under the central parts of the Mg II, at 2800 Å, as adopted in the creation of UV Fe II observational templates (e.g. Vestergaard & Wilkes, 2001). However, due to the overlap of the Fe II and Mg II there, it is much easier to measure the strong wings that Fe II forms at both the red and blue sides of Mg II. Proper modeling of the underlying Fe II may affect the conclusions about the dynamics of the Mg II emitting region (e.g. Jonić et al., 2016; Popović et al., 2019).
The contour plots for the ratio of red (2800-3090 Å) to blue (2650-2800 Å) wings of UV Fe II are shown in Figure 2. The plot of the ratio shows a complicated structure. We again see a trend similar to the one displayed in Figure 1. The same values of the ratio are found at low-density high ionization flux and high-density low ionization flux. At low ionization flux, the ratio again depends mostly on the density. However, in this case, we additionally observe the dependence on the turbulent velocity, as in the central parts the ratio is initially rising, and finally decreasing when the turbulent velocity crosses the value 50 km s. Maybe this is connected with the fact that the overall range in the measured ratio is much smaller for the red to the blue wing (typical range between 0.3 and 1.8) than in the case of the entire UV to optical range covering a broad range of magnitudes.
@@ 3.3 The effect of the UV Fe II broadening on the measurements of the outflow properties
In the analysis of their Sloan Digital Sky Survey (SDSS) sample, Kovačević-Dojčinović & Popović (2015) noticed a systematic averaged redshift in the UV Fe II to the optical Fe II. To check if such a shift can be caused just by the shape of the Fe II emission when broadened with the turbulence effect included, we used a specific Fe II template characterized by the ionization flux, , and the density, . We applied the Gaussian smearing to the template, and we determined the peak wavelength of the blue wing of UV Fe II. The peak is usually located close to Å , and it is dominated by the multiplets 62 and 63 (see Kovačević-Dojčinović & Popović 2015). However, when we model the entire Fe II emission of the blue wing we see that the position of the peaks moves with the adopted smear velocity. We plot this effect in Figure 3 simply in Å units, and in Figure 4 in velocity space. The shift corresponds to the apparent velocity from 100 km s up to 300 km s when the turbulence is not included, and the presence of the turbulence affects the apparent direction of the outflow and its amplitude. Apparent velocity shifts up to 1000 km s are possible even though no velocity is present in the model which corresponds in each case to the rest frame spectrum. In Kovačević-Dojčinović & Popović (2015) more than half of the sources show a kinematic offset above 1000 km s. We do not see such large shifts due to the Fe II smoothing but we experimented only with a single template.
3.4 Red-to-blue optical Fe II wings ratio
@@
This part of Fe II emission is important for accurate modelling of the H zone, as recently stressed by Popović et al. (2023). Proper modelling of the Fe II pseudo-continuum may strongly affect the conclusions about the dynamics of the BLR (like the presence or absence of the very broad line component to H, and the role of the intermediate line region), and the modelling of the quasar main sequence (Panda et al., 2018, 2019a, 2019b; Marziani et al., 2021).
@@
We show contour plots for the ratio of red (5150-5350 Å) to blue (4434-4684 Å) parts of optical Fe II in Figure 5. This ratio shows again the non-monotonic dependence of the ionization parameter but interestingly, almost no dependence on the density. The minimum value of around 1.0 is present at the ionization parameter cm s, increasing when we move both up and down in the ionization flux. The dependence on the turbulent velocity is marginally noticeable, particularly for this minimum value. Some traces of dependence on the density are only seen for very low ionization flux, below cm s, and only for very high densities. The value of the turbulent velocity affects the extension of this region.
Abundance (in ) | , | UV Fe II/optical Fe II | UV red/blue wings | Optical red/blue wings | R |
1 | 18,10 | 2.56 | 2.13 | 1.81 | 0.91 |
20.5,12 | 8.98 | 1.26 | 0.84 | 1.09 | |
5 | 18,10 | 2.28 | 2.05 | 1.99 | 1.27 |
20.5,12 | 5.63 | 1.21 | 0.75 | 2.49 | |
20 | 18,10 | 3.03 | 1.77 | 1.88 | 0.99 |
20.5,12 | 4.94 | 1.03 | 0.76 | 4.42 |
@@
When computing the Fe II emission with CLOUDY we also have, as a by-product, the intensities of other emission lines. The ratio of the optical Fe II emission to H plays an exceptional role in the EV1 of the quasar sequence (Boroson & Green, 1992; Sulentic et al., 2000; Shen & Ho, 2014; Marziani et al., 2018; Panda et al., 2018, 2019b; Du & Wang, 2019; Panda & Marziani, 2023). We estimated as the ratio of the integrated optical Fe II flux in the standard wavelength range 4434-4684 Å to the flux of H at 4861.32 Å. We present the corresponding contour maps in Figure 6.
@@
These plots show that most of the map area corresponds to the value of smaller than 1. Higher values, above 3, are only predicted for the intermediate values of the turbulent velocity (20 - 30 km s, and only for very high values of the ionization parameter () which supports the earlier findings by Panda et al. (2018) and Panda (2021). Smaller values of correspond to the intermediate A-type quasars. The basic classification of quasars into class A and B, introduced by Sulentic et al. (2000) was analogous to the older classification of Seyfert galaxies into Narrow Line Seyfert 1 (NLS1) class and other Seyfert 1 galaxies (Osterbrock & Pogge, 1985), only the limiting velocity was different (4000 km s vs. 2000 km s), which reflected the systematic difference in the mass of the central black hole. In both cases, one class (NLS1 and type A) of objects corresponded to relatively higher Eddington rate sources than the other class, which was reflected in line properties (intensities and shapes). Later, the specific ranges of were used to explore full 2-D classification (see e.g., Marziani et al., 2018). A-type quasars with were named Extreme Population A (xA) sources, and they were found to be applicable as standard candles for cosmology (Marziani & Sulentic, 2014; Dultzin et al., 2020; Marziani, 2023; Panda & Marziani, 2023). Most of the xA sources have the values of between 1 and 2 (Marziani et al., 2018) but Śniegowska et al. (2018) found one source with a large ratio. A few more such sources were reported by Shen et al. (2011) but as discussed in Śniegowska et al. (2018), those measurements were not of high quality. Moreover, the sources with 2 are reported in the newer catalogs (e.g. Negrete et al., 2018; Rakshit et al., 2020, 2021; Wu & Shen, 2022).
@@
We, thus, could conclude, from Figure 5, that xA sources require high-density BLR () and very high radiation flux. However, our templates were composed of solar metallicity. Some authors argued (e.g. Panda et al., 2019b; Śniegowska et al., 2021; Panda, 2021; Marziani et al., 2024) that these sources, in addition, require highly super-solar metallicity.
@@
We compared these predictions with the statistical distribution of the reported on the basis of SDSS quasar studies. We used two data sets: older massive fits to the SDSS quasars, based on data release 7 (DR 7), were reported by Shen et al. (2011). Newer results, based on DR 14, were reported by Rakshit et al. (2020). The results are based on the spectral decomposition of the Fe II emission integrated within the traditional range 4434-4684 Å (Boroson & Green, 1992). In both cases, we selected only the sources with reliable spectra decomposition by requiring that the relative error of the FWHM of H and the measurement are below 20%. The resulting histograms, shown in Figure 7, consistently show the peak of the values at about 0.6 in both samples. There are very few objects with values above 4.0 which we ignored for better clarity of the plot. Such values are present either in the region of higher density and high ionization flux, above cm s, or low ionization flux and low density (lower left corner of plots in the Figure 6). The first location seems more likely as it is smoothly connecting to even larger values of than the average value, and the observational histogram also smoothly extends to higher values. However, we should stress that the observational results are not based on the Fe II emission models presented in this paper. Results of both Shen et al. (2011) and Rakshit et al. (2020) are based on the use of the optical Fe II template from Boroson & Green (1992).
@@ 3.6 The effect of metallicity
To investigate the effect of metallicity on Fe II emission, we rebuild our models for abundances equal to 2, 5, 7, 10, and 20 times solar abundances, with a fixed microturbulence velocity of 20 km/s. We plot contours for the ratios estimated in the previous sections in Figures 8-11 for various abundances. The values of these ratios at two arbitrary positions are given in Table 2 to quantify their variations with abundances.
The UV Fe II to optical Fe II contours, shown in Figure 8, exhibit marginal changes with abundances. The ratio decreases with increasing abundance at (=(20.5,12), however at no such trend is seen at lower incident flux, and lower density (=(18,10). Verner et al. (2003) investigated the effect of varying abundance on the ratio of UV Fe II to optical Fe II and found that the ratio decreases from 8 to 5.6 when abundance increases from solar to 10 times solar abundance for and a microturbulence velocity of 1 km/s. They concluded that the optical Fe II emission increases more than the UV Fe II emission when the abundance increases which is supported by the findings from Panda et al. (2019a).
Figure 9 shows the contours for the ratio of UV wings, whereas Figure 10 shows those of optical wings. As seen from the Figures and Table 2, both ratios are essentially unaffected by the abundance.
The contours for R are plotted in Figure 11 which shows small variations with abundance. The ratio increases with increasing abundance at high ionization flux and high density (=(20.5,12), but no such trend is seen at (=(18,10).
@@ 3.7 Fe II templates
For convenient use in future spectral fitting, we created a new family of Fe II templates. They come from the simulations as described in Section 2, i.e. those are 2646 models covering the range of the local density, ionization flux and turbulent velocity. Output files from CLOUDY are usually large, and data fitting does not require full resolution available in these output files since most of AGN have broad Fe components. Therefore, we binned the output spectra using the 2 Å bins. We tested whether this does not lead to a considerable problem in data fitting by assuming that the AGN spectrum comes from the medium with the velocity dispersion 2000 km s and we compared the Fe II spectrum in the 2650 - 2850 Å range without any prior binning, and with a bin size of 1, 2 and 5 Å. The result is shown in Figure 12. Up to the bin size of 2 Å there is no net difference in the broadened spectra but the bin size of 5 Å seems to imply already a systematic difference, so we adopted the 2 Å bin size optimizing between the accuracy and the size of the file. The complete set of Fe II templates for solar metallicity is available through zenodo111https://zenodo.org/doi/10.5281/zenodo.10532690 and GitHub222https://github.com/Ashwani-88/Fe2_template links.
Source | RA | Dec | Redshift | log(L[erg/s]) | log ( [M]) | Reference | |
HE 0413-4031 | 1.39117 | 46.74 | 8.87 | 1.66 | Prince et al. (2023) | ||
SDSS RM102 | 0.86114 | 45.00 | 7.92 | 0.329 | Shen et al. (2019) |
The redshift to this source was recently firmly established from the narrow [OIII] line using the near-IR spectrum (Zajaček et al., 2023b).
3.8 Exemplary cases
We test how the new Fe II templates compare with the observed AGN spectra. Since we use the SED more appropriate for bright quasars to construct the Fe II templates, we selected two exemplary sources of this type namely, HE 0413-4031 and SDSS RM102. The basic properties of these sources are mentioned in Table 3. In the case of HE 0413-4031, only narrow UV band data are available while the spectrum of SDSS RM102 continuously covers the broad UV-to-optical band.

### v2

@@ 3.3 Red-to-blue UV Fe II wings ratio
In our next plot, we concentrate on the global properties of the UV part of Fe II emission which is important for the proper modeling of the Mg II line. The Fe II emission under the Mg II line is never close to zero, even right under the central parts of the Mg II, at 2800 Å, as adopted in the creation of UV Fe II observational templates (e.g. Vestergaard & Wilkes, 2001). However, due to the overlap of the Fe II and Mg II there, it is much easier to measure the strong wings that Fe II forms at both the red and blue sides of Mg II. Proper modeling of the underlying Fe II may affect the conclusions about the dynamics of the Mg II emitting region (e.g. Jonić et al., 2016; Popović et al., 2019).
The contour plots for the ratio of red (2800-3090 Å) to blue (2650-2800 Å) wings of UV Fe II, shown in Figure 8, exhibit a complicated structure. At low ionizing photon flux, the ratio depends mostly on the density. At higher ionizing photon flux the dependence is more complex. We also observe the dependence on the microturbulence velocity, as in the central parts the ratio is initially rising, and finally decreasing when the microturbulence velocity crosses the value 50 km/s. The measured value of the ratio (shown by the red dashed contour) for the test source is 2.02 which can only be found for a very narrow range of parameters for the microturbulence velocity of 20 km/s. The two wings represent different multiplets in templates of Kovačević-Dojčinović & Popović (2015), so they may, or may not, come from the same region (see Section 4).
@@ 3.4 The effect of the UV Fe II broadening on the measurements of the outflow properties
In the analysis of their Sloan Digital Sky Survey (SDSS) sample, Kovačević-Dojčinović & Popović (2015) noticed a systematic averaged redshift in the UV Fe II to the optical Fe II. To check if such a shift can be caused just by the shape of the Fe II emission when broadened with the microturbulence effect included, we used a specific Fe II template characterized by the ionizing photon flux, , and the density, . We applied a Gaussian smearing of 2000 km s-1 to the template and determined the peak wavelength of the blue wing of UV Fe II. The peak is usually located close to Å , and it is dominated by the multiplets 62 and 63 (see Kovačević-Dojčinović & Popović 2015). However, when we model the entire Fe II emission of the blue wing we see that the position of the peaks moves with the adopted smear velocity. We plot this effect simply in Å units (top panel), and in velocity space (bottom panel) in Figure 9. The shift corresponds to the apparent velocity from 100 km s-1 up to 300 km s-1 when the microturbulence is not included, and the presence of the microturbulence affects the apparent direction of the outflow and its amplitude. Apparent velocity shifts up to 1000 km s-1 are possible even though no velocity is present in the model which corresponds in each case to the rest-frame spectrum. In Kovačević-Dojčinović & Popović (2015) more than half of the sources show a kinematic offset above 1000 km s-1. We do not see such large shifts due to the Fe II smoothing but we experimented only with a single template.
3.5 UV-to-optical Fe II emission
We calculated the global ratio of the UV to optical broadband Fe II emission with the integration range specified in Section 2.2. The results are shown in Figure 10 using contour plots. The representative values in the central parts of the plot are in the range 6 to 10.
As already argued by Joly (1981) and Wills et al. (1985), a ratio of the order of 1 in the relative intensities of the optical to UV lines implies a column density of about to cm-2 to balance Fe II line optical depths and the Balmer opacity. In our computations, we use the value of cm-2 but we hardly recover the ratio measured for our reference object RM 102. We find that this ratio grows rapidly as the ionizing photon flux and cloud density increase. Panda et al. (2019a) and Sarkar et al. (2021) stressed the role of the microturbulence velocity in the Fe II intensity but the global ratio of the UV to optical is relatively unaffected.
The ratio shown in Figure 10 represents the expectations based on the assumption of a single cloud responsible for both UV and optical emission, which may not be the case in actual objects. We address this issue in the discussion.
3.6 Red-to-blue optical Fe II blends ratio
@@
This part of Fe II emission is important for accurate modeling of the H zone, as recently stressed by Popović et al. (2023). Proper modeling of the Fe II pseudo-continuum may strongly affect the conclusions about the dynamics of the BLR (like the presence or absence of the very broad line component to H, and the role of the intermediate line region), and the modeling of the quasar main sequence (Panda et al., 2018, 2019a, 2019b; Marziani et al., 2021b).
@@
We show contour plots for the ratio of red (5150-5350 Å) to blue (4434-4684 Å) parts of the optical Fe II in Figure 11. This ratio shows again the non-monotonic dependence of the ionizing photon flux but interestingly, almost no dependence on the density. The dependence on the microturbulence velocity is marginally noticeable. The measured value of this ratio in our test object is 0.68 (shown by red dashed contour) which could only be found for density cm-3, ionizing flux cm-2 s-1 and high microturbulence velocity.
@@
When computing the Fe II emission with CLOUDY we also have, as a by-product, the intensities of other emission lines. The ratio of the optical Fe II emission to H plays an exceptional role in the EV1 of the quasar sequence (Boroson & Green, 1992; Sulentic et al., 2000; Shen & Ho, 2014; Marziani et al., 2018; Panda et al., 2018, 2019b; Du & Wang, 2019; Panda & Marziani, 2023; Panda et al., 2024; Panda & Śniegowska, 2024). We estimated as the ratio of the integrated optical Fe II flux in the standard wavelength range 4434-4684 Å to the flux of H at 4861.32 Å. We present the corresponding contour maps in Figure 12.
@@
These plots show that most of the map area corresponds to the value of smaller than 1. Higher values, above 3, are only predicted for the intermediate values of the microturbulence velocity (20 - 30 km s, and only for very high values of the ionizing photon flux () which supports the earlier findings by Panda et al. (2018) and Panda (2021b). Smaller values of correspond to the early A and B type quasars. The basic classification of quasars into class A and B, introduced by Sulentic et al. (2000) was analogous to the older classification of Seyfert galaxies into Narrow Line Seyfert 1 (NLS1) class and other Seyfert 1 galaxies (Osterbrock & Pogge, 1985), only the limiting velocity was different (4000 km s-1 vs. 2000 km s-1), which reflected the systematic difference in the mass of the central black hole. Specific ranges of were used to explore full 2-D classification (see e.g., Marziani et al., 2018). A-type quasars with were named Extreme Population A (xA) sources, and they were found to be applicable as standard candles for cosmology (Marziani & Sulentic, 2014; Dultzin et al., 2020; Marziani, 2023; Panda & Marziani, 2023). The xA sources have values of between 1 and 2 (Marziani et al., 2018) but Śniegowska et al. (2018) found one source with a large ratio. A few more such sources were reported by Shen et al. (2011) but as discussed in Śniegowska et al. (2018), those measurements were not of high quality. Moreover, the sources with 2 are reported in the newer catalogs (e.g. Negrete et al., 2018; Rakshit et al., 2020, 2021; Wu & Shen, 2022; Panda et al., 2024).
Assuming a simple power-law shape for the continuum, we observed RFeII = 1.24 0.7 for RM 102. When we modeled the continuum as a combination of a disk and stellar contributions, the revised value is RFeII = 0.95 0.35. This decrease is attributed to the inclusion of the host galaxy, which can lead to potential misclassification of a source as xA, as demonstrated by Bon et al. (2020). However, both measurements are consistent within uncertainties and meet the criterion for xA sources (R). Additionally, Shen et al. (2019) reported an Eddington ratio of = 0.329 for this source. Both the Eddington ratio and RFeII measurements confirm that RM 102 belongs to the xA class of AGN.
@@
We noticed that the observed RFeII value for RM 102 falls within two distinct regions of the contour plots in Figure 12: one corresponding to low-density and low-ionizing flux and the other to high-density and high-ionizing flux. Previous studies on UV emission line ratios in xA sources (e.g. Negrete et al., 2012; Panda et al., 2019b) have favoured the possibility of a high-density ( (cm-3) = 12 - 13) BLR for these objects. We, thus, may conclude, from Figure 12, that xA sources require a high-density BLR ( (cm-3) = 12 - 13) and very high () ionizing flux. However, our templates were composed of solar metallicity. Some authors argued (e.g. Martínez-Aldama et al., 2018; Panda et al., 2019b; Śniegowska et al., 2021; Panda, 2021b; Marziani et al., 2024; Floris et al., 2024) that these sources, in addition, require highly super-solar metallicity.
@@
We compared these predictions with the statistical distribution of the reported based on SDSS quasar studies. We used two data sets: older massive fits to the SDSS quasars, based on data release 7 (DR 7), were reported by Shen et al. (2011). Newer results, based on DR 14, were reported by Rakshit et al. (2020). The results are based on the spectral decomposition of the Fe II emission integrated within the traditional range 4434-4684 Å (Boroson & Green, 1992). In both cases, we selected only the sources with reliable spectra decomposition by requiring that the relative error of the FWHM of H and the measurement are below 20%. The resulting histograms, shown in Figure 13, consistently show the peak of the values at about 0.6 in both samples. There are very few objects with values above 4.0 which we ignored for better clarity of the plot. Such values are present either in the region of higher density and high ionizing photon flux, above cm-2 s-1, or low ionizing photon flux and low density (lower left corner of plots in the Figure 12). The first location seems more likely as it is smoothly connecting to even larger values of than the average value, and the observational histogram also smoothly extends to higher values. However, we should stress that the observational results are not based on the Fe II emission models presented in this paper. Results of both Shen et al. (2011) and Rakshit et al. (2020) are based on the use of the optical Fe II template from Boroson & Green (1992).
Abundance (in ) | , | UV Fe II/optical Fe II | UV red/blue wings | Optical red/blue blends | RFeII |
1 | 18,10 | 4.96 | 1.59 | 1.80 | 1.20 |
20.5,12 | 11.82 | 1.23 | 1.03 | 0.60 | |
5 | 18,10 | 4.13 | 1.63 | 1.91 | 1.59 |
20.5,12 | 6.62 | 1.49 | 0.90 | 1.51 | |
10 | 18,10 | 4.17 | 1.64 | 1.91 | 1.47 |
20.5,12 | 5.33 | 1.56 | 0.86 | 2.04 |
@@ 3.8 The effect of metallicity
To investigate the effect of metallicity on Fe II emission, we rebuild our models for abundances equal to 2, 5, and 10 times solar abundances, with a fixed microturbulence velocity of 100 km/s. We plot contours for the ratios estimated in the previous sections in Figures 14-17 for various abundances. The values of these ratios at two arbitrary positions are given in Table 3 to quantify their variations with abundances.
The UV Fe II to optical Fe II contours, shown in Figure 14, exhibit marginal changes with abundances. The ratio decreases with increasing abundance at (=(20.5,12), however, no such trend is seen at lower incident flux, and lower density (=(18,10). Verner et al. (2003) investigated the effect of varying abundance on the ratio of UV Fe II to optical Fe II and found that the ratio decreases from 8 to 5.6 when abundance increases from solar to 10 times solar abundance for and a microturbulence velocity of 1 km/s. They concluded that the optical Fe II emission increases more than the UV Fe II emission when the abundance increases which is supported by the findings from Panda et al. (2019a).
Figure 15 shows the contours for the ratio of UV wings, whereas Figure 16 shows those of optical blends. As seen from the Figures and Table 3, both ratios are essentially unaffected by the abundance.
The contours for RFeII are plotted in Figure 17 which shows small variations with abundance. The ratio increases with increasing abundance at high ionizing photon flux and high density (=(20.5,12), but no such trend is seen at (=(18,10).
@@ 3.9 Fe II templates
For convenient use in future spectral fitting, we created a new family of Fe II templates. They come from the simulations as described in Section 2, i.e. those are 2646 models covering the range of the local density, ionizing photon flux and microturbulence. Output files from CLOUDY are usually large, and data fitting does not require full resolution available in these output files since most of AGN have broad iron components. Therefore, we binned the output spectra using the 2 Å bins. We tested whether this does not lead to a considerable problem in data fitting by assuming that the AGN spectrum comes from the medium with the velocity dispersion 2000 km s-1 and we compared the Fe II spectrum in the 2650 - 2850 Å range without any prior binning, and with a bin size of 1, 2 and 5 Å. The result is shown in Figure 18. Up to the bin size of 2 Å there is no net difference in the broadened spectra but the bin size of 5 Å seems to imply already a systematic difference, so we adopted the 2 Å bin size optimizing between the accuracy and the size of the file.
Additionally, we include templates for the inward and outward Fe II emission, which are split into 1000 logarithmic wavelength bins, each 584 km s-1 wide, between 1000 and 7000 Å. The complete set of Fe II templates for solar metallicity is available through zenodo222https://zenodo.org/doi/10.5281/zenodo.10532690 and GitHub333https://github.com/Ashwani-88/Fe2_template links. Each template corresponds to a single set of cloud parameters such as local density (nH), ionizing photon flux (), and microturbulence (). In the next Section, we compare our several Fe II templates based on single cloud parameters with observed Fe II spectra to comprehend the physical properties of the BLR cloud emitting Fe II emission. However, it is possible that a combination of clouds having different physical conditions can be responsible for the observed Fe II spectra as well as the other strong emission lines. Our Fe II templates can be combined using any multi-cloud model such as the Locally Optimized Cloud (LOC) model suggested by Baldwin et al. (1995) to compare the observed Fe II spectra of quasars.
Reverberation mapping studies indicate that the broad-line region (BLR) is radially extended and likely exhibits vertical stratification, along with a range of local densities due to thermal instabilities and the cloud formation process. Our publicly available set of Fe II templates is thus envisioned as a good starting point for further modelling, using LOC or other BLR models which include stratification. However, such modelling is beyond the scope of our current objective.
3.10 Testing the templates against the complete spectra of the two test objects
We test how the new Fe II templates compare with the observed AGN spectra instead of using just integrated spectral bands. In the case of HE 0413-4031, only narrow UV band data are available while the spectrum of SDSS RM102 continuously covers the broad UV-to-optical band.